Meeting 2022 08 10
GMn GEM commissioning Meetings: WEDNESDAY 10:30am - 11:30am
Link to all previous meetings: GMn GEM Commissioning Meeting
Please use sbs_gems mailing list for plots/communications and HALOG entries relevant to GEMs!
Near term work plan
- The second fiber patch panel and the fiber cable were installed
- All the 3 VME MPD crates are on the network now and we are carrying out low level tests now.
- Remote controlled power strips in the SBS GEM bunker are also online and working.
- SBS gas system - Jack
- Manpower for August
- See installation plan for SBS here: plan
- [photos from the hall https://hallaweb.jlab.org/wiki/index.php/File:Installation_Update_22June2022.pdf]
- Installation of HV dividers on BB: work plan
- Thoughts on special run requests during GEn:
- Luminosity studies (Nilanga)
- Full readout studies for negative pulse studies and etc (Sean)
--- Current Status of SBS Data/Analyzer ---
- Took cosmics and saw signal after a latency scan.
- For final latency settings it is easier to tune this using the Analyzer/SBS Offline using the gem_basic cfg
- For raw data/GEM viewer:
- Looks okay so far. Looks like some disconnected APVs/cables (eaiser to see the in raw viewer)
- Mapping looks okay
- Next step would be to get pedestal and apply Zero Suppression and Common Mode
- Then re-take cosmics with these turned on (Long run)
- For Analyzer/SBS Offline:
- Need to finish/check mapping (a bit more to this one)
- Need to run pedestal and common mode
- Need to take run with these ON
- Need to do alignment
- Need to re-take data
- Then analyze --> May see issues here with mapping, alignment, etc. etc. etc.
Update on INFN GEM layers for GEn-RP
- Ran up a new divider on the front layer GEM of the inline stack. Verified temperature and V/I consistent with divider run-up with no GEM.
- Currently in Hall. Plan to install new dividers on upstream-most layer while we can easily access. Will need sliding platform to remove the layer to install the new dividers on the second layer.
Update on UVa GEM layers for GEn-RP
- Gas shut-off valve configuration for outputs were changed as per the advice from Nilanga and Huong. Now we have two shut-off valves for the two separate outputs instead of T-ing the outputs and putting a single shut-off valve. This work is only partially completed as we had to order more shut-off valves for this. In the meantime, we disconnected the unfinished GEMs from the T connectors.
- We successfully completed out very first HV test under Nitrogen in Hall! See the log entry: SBS Inline UVA GEM first HV test under nitrogen in the Hall
- Majority of the cable connections/clock phase are fine for the 6 UVA layers. We still have to re-test a few MPDs/APVs today and hopefully we will be done within today itself.
- What's left after that will be take pedestal data.
Update on BB GEMs
- All BB HV now immediately shuts off if the gas mixture is out of an acceptable range.
- HiVolta supply channels are all forced to be on or off together.
- BB XY layer investigation